# VMarker-Pro: Probabilistic 3D Human Mesh Estimation from Virtual Markers

### v3

@@ 3.2.1 VMarker
3D VM estimation. We train a neural network to estimate a 3D heatmap $\mathbf{H}=[\mathbf{H}_{1},\,...,\,\mathbf{H}_{K}]\in\mathbb{R}^{K\times D\times H\times W}$ from an image. The heatmap encodes per-voxel likelihood of each marker. There are $D\times H\times W$ voxels in total which are used to discretize the 3D space. The 3D position $\mathbf{P}_{z}\in\mathbb{R}^{3}$ of each marker is computed as the center of mass of the corresponding heatmap $\mathbf{H}_{z}$ [18] as follows:
$\mathbf{P}_{z}=\sum_{d=1}^{D}\sum_{h=1}^{H}\sum_{w=1}^{W}(d,h,w)\cdot\mathbf{H}_{z}(d,h,w).$ (3)
The positions of all markers are represented as $\mathbf{P}=[\mathbf{P}_{1},\mathbf{P}_{2},\cdots,\mathbf{P}_{K}]$.
3D mesh estimation. Ideally, if we have accurate estimates for all virtual markers $\mathbf{P}$, then we can recover the complete mesh by simply multiplying $\mathbf{P}$ with a fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$ with sufficient accuracy as validated in Table I. However, in practice, some markers may have large estimation errors because they may be occluded in the monocular setting. Note that this happens frequently. For example, the markers in the back will be occluded when a person is facing the camera. As a result, inaccurate markers positions may bring large errors to the final mesh if we directly multiply them with the fixed matrix $\widetilde{\mathbf{A}}^{sym}$.
Our solution is to rely more on those accurately detected markers. To that end, we propose to update the coefficient matrix based on the estimation confidence scores $\mathbf{C}$ of the markers. In practice, we simply take the heatmap score at the estimated positions of each marker obtained by interpolation. For convenience, we denote this process as $\mathbf{C}_{z}=\mathbf{H}_{z}(\mathbf{P}_{z})$. We feed the confidence scores to an updating network which is a single fully-connected (FC) layer to obtain the coefficient matrix $\mathbf{A}$. Then the mesh is reconstructed by linear multiplication $\mathbf{M}=\mathbf{P}\mathbf{A}$.
3.2.2 Training
@@
$\mathcal{L}_{vm}=\lambda_{3d}\mathcal{L}_{3d}+\lambda_{c}\mathcal{L}_{conf}+\lambda_{m}\mathcal{L}_{mesh}.$ (4)
Virtual marker loss. We define $\mathcal{L}_{3d}$ as the $L_{1}$ distance between the predicted 3D virtual markers $\mathbf{P}$ and the GT $\hat{\mathbf{P}}$ as follows:
$\mathcal{L}_{3d}=\|\mathbf{P}-\hat{\mathbf{P}}\|_{1}.$ (5)
Note that it is easy to get GT markers $\hat{\mathbf{P}}$ from GT meshes as stated in Section 3.1 without additional manual annotations.
Confidence loss. We also require that the 3D heatmaps have reasonable shapes, therefore, the heatmap score at the voxel containing the GT marker position $\hat{\mathbf{P}}_{z}$ should have the maximum value as in the previous work [72]:
$\mathcal{L}_{conf}=-\sum_{z=1}^{K}log(\mathbf{H}_{z}(\hat{\mathbf{P}}_{z})).$ (6)
@@
$\mathcal{L}_{mesh}=\mathcal{L}_{vertex}+\mathcal{L}_{pose}+\mathcal{L}_{normal}+\lambda_{e}\mathcal{L}_{edge}.$ (7)
  • Vertex coordinate loss. We adopt $L_{1}$ loss between predicted 3D mesh coordinates $\mathbf{M}$ with GT mesh $\hat{\mathbf{M}}$ as:
    $\mathcal{L}_{vertex}=\|\mathbf{M}-\hat{\mathbf{M}}\|_{1}.$ (8)
  • Pose loss. We use $L_{1}$ loss between the 3D landmark joints regressed from mesh $\mathbf{M}\mathcal{J}$ and the GT joints $\hat{\mathbf{M}}\mathcal{J}$ as:
    $\mathcal{L}_{pose}=\|\mathbf{M}\mathcal{J}-\hat{\mathbf{M}}\mathcal{J}\|_{1},$ (9)
    where $\mathcal{J}\in\mathbb{R}^{M\times J}$ is a pre-defined joint regression matrix in SMPL model [29] and $J$ is the number of joints.
@@
    $\mathcal{L}_{normal}=\sum_{f}\sum_{\{i,j\}\subset f}\left|\left<\frac{\mathbf{M}_{i}-\mathbf{M}_{j}}{\|\mathbf{M}_{i}-\mathbf{M}_{j}\|_{2}},\hat{\mathbf{n}}_{f}\right>\right|,$ (10)
    $\mathcal{L}_{edge}=\sum_{f}\sum_{\{i,j\}\subset f}\left|\|\mathbf{M}_{i}-\mathbf{M}_{j}\|_{2}-\|\hat{\mathbf{M}}_{i}-\hat{\mathbf{M}}_{j}\|_{2}\right|.$
@@
    where $f$ and $\hat{\mathbf{n}}_{f}$ denote a triangle face in the mesh and its GT unit normal vector, respectively. $\mathbf{M}_{i}$ denote the $i^{th}$ vertex of $\mathbf{M}$. $\hat{}$ denotes GT.
3.3 Probabilistic mesh estimation framework
3.3.1 VMarker-Pro
As shown in Figure 2, virtual markers are often self-occluded, thus there may exist multiple viable solutions. An incorrect estimation may introduce errors into the final mesh estimate. Therefore, we propose a probabilistic framework for modeling the estimation of 3D VMs, as illustrated in Figure 3 (bottom). By leveraging the powerful modeling capabilities of diffusion models [26, 74] on data distributions, we generate multiple feasible 3D VM estimates given an image condition. Specifically, we model the estimation of 3D VMs as a conditional reverse denoising process that starts with Gaussian noise and progressively denoises it to obtain the 3D VM using a denoiser network. In the following, we detail the problem formulation of the 3D VM estimation using diffusion models, including the definition of the forward diffusion process and the reverse denoising process. Then, we introduce the architecture design of the denoiser network.
Problem formulation. We formulate the 3D VM estimation as a reverse denoising process [26], as shown in the inference workflow at the bottom of Figure 3. Starting with a Gaussian noise $\mathcal{N}(\mathbf{0},\mathbf{I})$, we design a denoiser network to iteratively estimate the clean 3D VM. To achieve this, the framework includes two core processes: a forward diffusion process and a reverse denoising process. (1) The forward diffusion process disturbs a clean 3D VM, i.e.  the ground-truth (GT) 3D VM $\hat{\mathbf{P}}$ from the data distribution, by incrementally adding noise to generate a series of noisy 3D VMs and ultimately reaches a Gaussian prior. (2) Then the reverse denoising process iteratively denoises the Gaussian noise into clean data in a step-by-step manner. To guide the denoising process, we condition it on the 2D image cues $\mathbf{F}$ and the estimated 2D VMs $\mathbf{P}^{2d}\in\mathbb{R}^{2\times K}$, which is the first two dimensions of the estimated 3D VMs $\mathbf{P}$ by VMarker. By formulating the estimation as a reverse denoising process, we leverage the generative capabilities of diffusion models, framing the task as a probabilistic estimation naturally.
Forward diffusion process. Starting from a clean 3D VM $\hat{\mathbf{P}}$ drawn from the GT data distribution, also denoted as $\hat{\mathbf{P}}(0)$, we construct a time-dependent diffusion process by incrementally adding noise to $\hat{\mathbf{P}}(0)$ to generate a series of noisy 3D VMs $\{\hat{\mathbf{P}}(t)\}_{t=0}^{T}$, indexed by a discrete timestep $t\in\{0,1,...,T\}$, where $T$ denotes the total number of timesteps. Formally, we define a noisy 3D VM $\hat{\mathbf{P}}(t)$ at timestep $t$ as:
$q(\hat{\mathbf{P}}(1:T)|\hat{\mathbf{P}}(0))$ $:=\prod_{t=1}^{T}q(\hat{\mathbf{P}}(t)|\hat{\mathbf{P}}(t-1)),$ (11)
$q(\hat{\mathbf{P}}(t)|\hat{\mathbf{P}}(t-1))$ $:=\mathcal{N}(\hat{\mathbf{P}}(t);\sqrt{1-\beta_{t}}\hat{\mathbf{P}}(t-1),\beta_{t}\mathbf{I}),$ (12)
where $\{\beta_{t}\}_{t=1}^{T}$ is a variance schedule [26]. Leveraging the additive property of independent Gaussian distributions and reparameterization techniques [75], we can obtain the perturbation $\hat{\mathbf{P}}(t)$ from a clean 3D VM $\hat{\mathbf{P}}(0)$ directly:
$q(\hat{\mathbf{P}}(t)|\hat{\mathbf{P}}(0)):=\sqrt{\overline{\alpha}_{t}}\hat{\mathbf{P}}(0)+\sqrt{1-\overline{\alpha}_{t}}\hat{\bm{\epsilon}},$ (13)
where $\alpha_{t}:=1-\beta_{t}$, $\overline{\alpha_{t}}:=\prod_{u=1}^{t}\alpha_{u}$, and $\hat{\bm{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ as a Gaussian noise. Following previous work [26], we disturb any clean 3D VM $\hat{\mathbf{P}}(0)$ to a Gaussian distribution $\hat{\mathbf{P}}(T)\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
Reverse denoising process. Starting from a Gaussian distribution, denoted as $\mathbf{P}(T)\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, we reverse the diffusion process by incrementally denoising it to get a clean 3D VM $\mathbf{P}(0)$. We adopt Denoising Diffusion Implicit Models (DDIM) [60] to accelerate the reverse denoising process with fewer timesteps while keeping the denoising quality. Formally, following DDIM [60], we define a timestep subset $\tau\subset\{0,1,...,T\}$, and the reverse process is defined as:
$\begin{split}\mathbf{P}({\tau_{i-1}})&=\sqrt{\overline{\alpha}_{\tau_{i-1}}}(\frac{\mathbf{P}({\tau_{i}})-\sqrt{1-\overline{\alpha}_{\tau_{i}}}\bm{\epsilon}_{\tau_{i}}}{\sqrt{\overline{\alpha}_{\tau_{i}}}})\\ &+\sqrt{1-\overline{\alpha}_{\tau_{i-1}}-\sigma_{\tau_{i}}^{2}}\bm{\epsilon}_{\tau_{i}}+\sigma_{\tau_{i}}\hat{\bm{\epsilon}}_{\tau_{i}},\end{split}$ (14)
where $\tau_{i},\tau_{i-1}$ are the adjacent timesteps in the subset $\tau$. $\sigma_{\tau_{i}}(\eta)=\eta\sqrt{(1-\overline{\alpha}_{\tau_{i-1}})/({1-\overline{\alpha}_{\tau_{i}}})}\sqrt{1-\overline{\alpha}_{\tau_{i}}/\overline{\alpha}_{\tau_{i-1}}}$, where $\eta\in\mathbb{R}_{\geq 0}$ is a hyperparameter to control the generation process. $\hat{\bm{\epsilon}}_{\tau_{i}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is standard Gaussian noise.
In order to simulate Eq. 14 to perform the denoising process, we need to know $\bm{\epsilon}_{\tau_{i}}$ for all timestep $\tau_{i}$. Therefore, we train a time-dependent denoiser network to estimate it as shown at the bottom block of Figure 3. Instead of directly predicting $\bm{\epsilon}_{\tau_{i}}$ as the original formulation by Ho et al.  [26], following previous work [76, 68], we predict the clean signal itself, i.e.  $\mathbf{P}(0)$, and obtain the noise $\bm{\epsilon}_{\tau_{i}}$ through:
$\bm{\epsilon}_{\tau_{i}}=\frac{\mathbf{P}({\tau_{i}})-\sqrt{\overline{\alpha}_{{\tau_{i}}}}\mathbf{P}(0)}{\sqrt{1-\overline{\alpha}_{\tau_{i}}}}.$ (15)
By combining Eq. 15 and Eq. 14, we can get the denoised $\mathbf{P}({\tau_{i-1}})$ as follows:
$\mathbf{P}({\tau_{i-1}})$ $=\sqrt{\overline{\alpha}_{\tau_{i-1}}}\mathbf{P}(0)$ (16)
$+\sqrt{1-\overline{\alpha}_{\tau_{i-1}}-\sigma_{\tau_{i}}^{2}}\frac{\mathbf{P}({\tau_{i}})-\sqrt{\overline{\alpha}_{{\tau_{i}}}}\mathbf{P}(0)}{\sqrt{1-\overline{\alpha}_{\tau_{i}}}}$
$+\sigma_{\tau_{i}}\hat{\bm{\epsilon}}_{\tau_{i}}.$
Denoiser architecture design. Figure 4 presents an overview of the time-dependent denoiser network architecture. The denoiser aims to estimate the clean 3D VM $\mathbf{P}(0)$ given a noisy 3D VM $\hat{\mathbf{P}}(t)$ at timestep $t$ as input. To guide the prediction, we concatenate $\hat{\mathbf{P}}(t)$ with the estimated 2D VM $\mathbf{P}^{2d}$, and then transform them to feature $\mathbf{F}^{vm}\in\mathbb{R}^{K\times C}$ by a multilayer perceptron (MLP), where $C$ denotes the number of feature channel. We then deploy a network with $B$ blocks to process the VM features $\mathbf{F}^{vm}$. Each block is composed of two Locally Connected Network (LCN) units [77, 78] sequentially. Within each unit, an LCN layer [77, 78] is followed by Group Normalization (GN) [79], a SiLU activation function [80], and Dropout [81] for regularization. The LCN layer, as proposed in [77, 78], is akin to an enhanced version of MLP, offering more effective feature processing across virtual markers.
To improve the alignment of the denoised 3D VM with 2D image cues, we extract 2D image features $\mathbf{F}\in\mathbb{R}^{H_{b}\times W_{b}\times C_{b}}$ using the CNN backbone [18]. These features are sampled according to the estimated 2D VMs $\mathbf{P}^{2d}$ and subsequently transformed into a new feature space $\mathbf{F}^{2d}\in\mathbb{R}^{K\times C}$ through a linear layer, as illustrated in Figure 4. We then fuse $\mathbf{F}^{2d}$ into the network by Cross-Attention [82]. To achieve this, we slightly modify the LCN layer by incorporating an attention module [82].
In addition, we process the current timestep $t$ using an MLP to be $\mathbf{F}^{t}\in\mathbb{R}^{1\times C}$ and add it to the processed VM features $\mathbf{F}^{vm}$ to make the denoiser time-dependent following common practice [74, 53]. Finally, an FC layer is deployed to estimate the clean signal $\mathbf{P}(0)$. Please refer to the supplementary for more details.
3.3.2 Training
The whole framework is trained in an end-to-end manner. The overall loss function is defined as:
$\mathcal{L}_{vmpro}=\mathcal{L}_{vm}+\lambda_{diffusion}\mathcal{L}_{diffusion},$ (17)
where $\mathcal{L}_{vm}$ is the loss function for VMarker, as defined in Eq. 4. Following common practice in training diffusion models [26, 68], we randomly sample a timestep $t\in\{0,1,...,T\}$ to get the noisy 3D VM $\hat{\mathbf{P}}(t)$ according to Eq. 13 and deploy $\mathcal{L}_{diffusion}$ as the $L_{1}$ distance between the predicted clean 3D VM $\mathbf{P}(0)$ and the GT $\hat{\mathbf{P}}(0)$:
$\mathcal{L}_{diffusion}=\mathbb{E}_{t,\hat{\mathbf{P}}(0),\hat{\bm{\epsilon}}}[\|\mathbf{P}(0)-\hat{\mathbf{P}}(0)\|_{1}].$ (18)
3.3.3 Inference
As shown in Figure 3, at inference, the denoiser iteratively removes noise from a Gaussian noise $\mathbf{P}(T)\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ using DDIM [60] according to Eq. 16, yielding a denoised 3D VM $\mathbf{P}(0)$. Finally, we obtain a mesh estimate $\mathbf{M}=\mathbf{P}(0)\mathbf{A}$. By sampling $S$ Gaussian noises, we can obtain $S$ feasible mesh estimates.
Figure 4: Architecture of the denoiser network, which predicts the clean 3D VM $\mathbf{P}(0)$ from the noisy 3D VM $\hat{\mathbf{P}}(t)$. To guide the direction of denoising, we additionally input 2D features $\mathbf{F}$, the current timestep $t$, and the 2D VM $\mathbf{P}^{2d}$ predicted by VMarker. We concatenate $\mathbf{P}^{2d}$ with the noisy 3D VM $\hat{\mathbf{P}}(t)$ to facilitate training and fuse the sampled 2D features $\mathbf{F}$ through a Cross-Attention mechanism [82]. We utilize a network comprising $B$ blocks to process the features, where each block contains two Locally Connected Network (LCN) units [77, 78]. Each LCN units consists of one LCN layer [77, 78], followed by Group Normalization (GN) [79], SiLU activation [80], and Dropout [81]. Finally, an FC layer is used to predict the clean 3D VM $\mathbf{P}(0)$.
@@ 4 Experiments
TABLE II: Comparison to the deterministic state-of-the-arts on H3.6M [13] and 3DPW [14] datasets. means using temporal cues. The methods are not strictly comparable because they may have different backbones and training datasets. We provide the numbers only to show proof-of-concept results.
Method Venue Intermediate H3.6M [13] 3DPW [14]
@@
VMarker (Ours) [28] CVPR’23 Virtual marker 58.0 47.3 32.0 77.9 67.5 41.3
TABLE III: Comparison to the probabilistic state-of-the-arts on H3.6M [13] and 3DPW [14] datasets. The methods are not strictly comparable because they may have different backbones and training datasets. We provide the numbers only to show proof-of-concept results.
Method Hypothesis H3.6M [13] 3DPW [14]
[…]
Biggs et al.  [41] NeurIPS’20 10 - 59.2 42.2 - 79.4 56.6
25 - 58.2 42.2 - 75.8 55.6
Sengupta et al.  [56] ICCV’21 25 - - - - 75.1 47.0
ProHMR [44] ICCV’21 10 - - - - 88.9 55.0
25 - - - - 85.1 52.1
100 - - - - 80.1 48.1
200 - - - - 77.9 46.5
HuManiFlow [43] CVPR’23 10 - - - - 75.6 47.9
25 - - - - 71.9 44.5
100 - - - - 65.1 39.9
200 - - - - 64.5 38.8
HMDiff [45] ICCV’23 25 - 49.3 32.4 82.4 72.7 44.5
10 55.8 45.4 31.3 77.3 66.6 42.5
25 52.9 42.9 29.8 72.9 63.2 40.0
100 49.2 39.9 27.6 67.4 58.2 36.4
VMarker-Pro (Ours) 200 47.7 38.6 26.6 64.7 56.2 35.0

### v4

@@ 3.2 Mesh estimation framework
3D marker estimation. We train a neural network to estimate a 3D heatmap $\hat{\mathbf{H}}=[\hat{\mathbf{H}}_{1},\,...,\,\hat{\mathbf{H}}_{K}]\in\mathbb{R}^{K\times D\times H\times W}$ from an image. The heatmap encodes per-voxel likelihood of each marker. There are $D\times H\times W$ voxels in total which are used to discretize the 3D space. The 3D position $\hat{\mathbf{P}}_{z}\in\mathbb{R}^{3}$ of each marker is computed as the center of mass of the corresponding heatmap $\hat{\mathbf{H}}_{z}$ [47] as follows:
$\small\hat{\mathbf{P}}_{z}=\sum_{d=1}^{D}\sum_{h=1}^{H}\sum_{w=1}^{W}(d,h,w)\cdot\hat{\mathbf{H}}_{z}(d,h,w).$ (3)
The positions of all markers are represented as $\hat{\mathbf{P}}=[\hat{\mathbf{P}}_{1},\hat{\mathbf{P}}_{2},\cdots,\hat{\mathbf{P}}_{K}]$.
Interpolation. Ideally, if we have accurate estimates for all virtual markers $\hat{\mathbf{P}}$, then we can recover the complete mesh by simply multiplying $\hat{\mathbf{P}}$ with a fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$ with sufficient accuracy as validated in Table 1. However, in practice, some markers may have large estimation errors because they may be occluded in the monocular setting. Note that this happens frequently. For example, the markers in the back will be occluded when a person is facing the camera. As a result, inaccurate markers positions may bring large errors to the final mesh if we directly multiply them with the fixed matrix $\widetilde{\mathbf{A}}^{sym}$.
Our solution is to rely more on those accurately detected markers. To that end, we propose to update the coefficient matrix based on the estimation confidence scores of the markers. In practice, we simply take the heatmap score at the estimated positions of each marker, i.e. $\hat{\mathbf{H}}_{z}(\hat{\mathbf{P}}_{z})$, and feed them to a single fully-connected layer to obtain the coefficient matrix $\hat{\mathbf{A}}$. Then the mesh is reconstructed by $\hat{\mathbf{M}}=\hat{\mathbf{P}}\hat{\mathbf{A}}$.
3.3 Training
@@
$\mathcal{L}=\lambda_{vm}\mathcal{L}_{vm}+\lambda_{c}\mathcal{L}_{conf}+\lambda_{m}\mathcal{L}_{mesh}.$ (4)
Virtual marker loss. We define $\mathcal{L}_{vm}$ as the $L_{1}$ distance between the predicted 3D virtual markers $\hat{\mathbf{P}}$ and the GT $\hat{\mathbf{P}}^{*}$ as follows:
$\mathcal{L}_{vm}=\|\hat{\mathbf{P}}-\hat{\mathbf{P}}^{*}\|_{1}.$ (5)
Note that it is easy to get GT markers $\hat{\mathbf{P}}^{*}$ from GT meshes as stated in Section 3.1 without additional manual annotations.
Confidence loss. We also require that the 3D heatmaps have reasonable shapes, therefore, the heatmap score at the voxel containing the GT marker position $\hat{\mathbf{P}}_{z}^{*}$ should have the maximum value as in the previous work [17]:
$\mathcal{L}_{conf}=-\sum_{z=1}^{K}log(\hat{\mathbf{H}}_{z}(\hat{\mathbf{P}}_{z}^{*})).$ (6)
@@
$\small\mathcal{L}_{mesh}=\mathcal{L}_{vertex}+\mathcal{L}_{pose}+\mathcal{L}_{normal}+\lambda_{e}\mathcal{L}_{edge}.$ (7)
  • Vertex coordinate loss. We adopt $L_{1}$ loss between predicted 3D mesh coordinates $\hat{\mathbf{M}}$ with GT mesh $\hat{\mathbf{M}}^{*}$ as:
    $\mathcal{L}_{vertex}=\|\hat{\mathbf{M}}-\hat{\mathbf{M}}^{*}\|_{1}.$ (8)
  • Pose loss. We use $L_{1}$ loss between the 3D landmark joints regressed from mesh $\hat{\mathbf{M}}\mathcal{J}$ and the GT joints $\hat{\mathbf{J}}^{*}$ as:
    $\mathcal{L}_{pose}=\|\hat{\mathbf{M}}\mathcal{J}-\hat{\mathbf{J}}^{*}\|_{1},$ (9)
    where $\mathcal{J}\in\mathbb{R}^{M\times J}$ is a pre-defined joint regression matrix in SMPL model [2].
@@
    $\mathcal{L}_{normal}=\sum_{f}\sum_{\{i,j\}\subset f}\left|\left<\frac{\hat{\mathbf{M}}_{i}-\hat{\mathbf{M}}_{j}}{\|\hat{\mathbf{M}}_{i}-\hat{\mathbf{M}}_{j}\|_{2}},\hat{\mathbf{n}}_{f}^{*}\right>\right|,$ (10)
    $\mathcal{L}_{edge}=\sum_{f}\sum_{\{i,j\}\subset f}\left|\|\hat{\mathbf{M}}_{i}-\hat{\mathbf{M}}_{j}\|_{2}-\|\hat{\mathbf{M}}_{i}^{*}-\hat{\mathbf{M}}_{j}^{*}\|_{2}\right|.$
@@
    where $f$ and $\hat{\mathbf{n}}_{f}^{*}$ denote a triangle face in the mesh and its GT unit normal vector, respectively. $\hat{\mathbf{M}}_{i}$ denote the $i^{th}$ vertex of $\hat{\mathbf{M}}$. denotes GT.
@@ 4 Experiments
Method Intermediate H3.6M 3DPW
@@
Ours Virtual marker 58.0 47.3 32.0 77.9 67.5 41.3
Table 2: Comparison to the state-of-the-arts on H3.6M [16] and 3DPW [54] datasets. means using temporal cues. The methods are not strictly comparable because they may have different backbones and training datasets. We provide the numbers only to show proof-of-concept results.